OpenStudy (anonymous):

an arhitect desings a rectangular flower garden such that the width is exactly two-thirds of he lenght. If 280 feet of antique picket fencing are to be used to enclose the garden, find the dimensions of the garden. (find lenght an width)

OpenStudy (anonymous):

 Let L = length, and W = width. We know the perimeter of this rectangle is 2L + 2W, and we know that W = (2/3)L So the perimeter of the rectangle is 2L + 2(2/3)L. We also know that the perimeter of the garden is 280. So, \[2L + 2(2/3)L = 280\] Solving for L gives us L = 84 Now we can use 2L + 2W = 280 to find the width by plugging in our value of L, \[2(84) + 2W = 280\] W = 56

OpenStudy (anonymous):

Also, sorry but the rectangle is not drawn to scale. xD.

OpenStudy (anonymous):

hehe, thank you for your quick answer! i really appreciate it